a) A normal 20-sided die has each of the numbers 1 through to 20 written on its faces, and has an equal chance of landing on any one of these numbers. If it is rolled 7 times, what is the probability that a 3 turns up exactly 5 times?
Answer #1

P(turning up 3) = 1/20 = 0.05

n = 7

P(X = x) = 7Cx * 0.05x * (1 - 0.05)7-x

P(X = 5) = 7C5 * 0.055 * 0.952 = 0.0000059 (ans)